OpenStudy (anonymous):

How many grams of iron metal do you expect to be produced when 245 grams of an 80.5 percent by mass iron (II) nitrate solution reacts with excess aluminum metal? Show all of the work needed to solve this problem. 2 Al (s) + 3 Fe(NO3)2 (aq) yields 3 Fe (s) + 2 Al(NO3)3 (aq)

OpenStudy (anonymous):

So far I have 245 g Fe(NO3)2 x (80.5 g Fe(NO3)2 solute/ 100 g Fe(NO3)2) x ( 1 mol Fe(NO3)2/ 179.85 g Fe(NO3)2) x (1 mol Al/2 mol Fe(NO3)2) x (26.98 g Al/1 mol Al)

OpenStudy (anonymous):

I don't know what to do from here please help

OpenStudy (australopithecus):

2 Al (s) + 3 Fe(NO3)2 (aq) -> 3 Fe (s) + 2 Al(NO3)3 (aq) 0.8*245 grams = Grams of iron (II) nitrate Next convert to moles moles = gram/molecular mass molecular mass can be determined from the periodic table just look up a youtube video if you dont know how to find it. From your formula for every 3 atoms of Fe(NO3)2 reacted 3 atoms of Elemental Iron are produced. So divide the moles by 3, then multiply by 3 to receive the amount of moles of elemental Iron, this gives you the moles of elemental iron produced. Then convert from moles to grams using the formula provided above Alternative way you could solve this is simply by using gravimetric factor (look it up) assuming the reaction goes to completion and conservation of mass you could just find the amount of iron in 0.8*245 grams of Iron (II) nitrate in one simple calculation

OpenStudy (anonymous):

can someone help me through the steps?? please I need to do this or ill get kicked out of my course :(

OpenStudy (abmon98):

Mass:80/100*245=197.225 Molar mass of iron natrate:179.8 Number of moles:Mass/Molar Mass Number of moles=197.225/179.8 Number of moles=1.096 Ratio of iron metal to iron nitrate is 3:3 so they share the same number of moles 1.096*55.8=61.2grams

OpenStudy (anonymous):

I got 61.435?

OpenStudy (anonymous):

0.805 x 245g = grams of iron II nitrate 197.26 = grams of iron II nitrate moles = grams / molecular mass moles = 197.26/ 179.845 moles = 1.10 1.10 moles Iron grams = moles x molecular mass grams = 1.10 x 55.85 grams = 61.435

OpenStudy (abmon98):

thats right

OpenStudy (anonymous):

okay cool thanks(:

OpenStudy (abmon98):

your welcome :)